OpenStudy (anonymous):

Calculate the average speed for each of the following. The only one I'm having trouble is this one: 292 1/2 km in 4 1/2 h. __________-

OpenStudy (anonymous):

just divide 292.5/4.5 and you get 65km/h

OpenStudy (anonymous):

Oh thanks :D Btw would you know how to find the missing value for this question? 2/15 / [ ] = 2/3

OpenStudy (anonymous):

Wouldn't it just be 2/15/5=2/3. Because 15/5 is 3. and 2/3 is 2/3. So 2/15/5 is 2/3.

OpenStudy (anonymous):

Wouldn't that be 2/45?

OpenStudy (anonymous):

I mean 2/75?